OpenStudy (secret-ninja):

Medals. Can someone walk me through how to do this? A system of equations is shown below: 8x + 5y = 9 3x + 2y = 4 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (secret-ninja):

@amistre64 @StudyGurl14 @KamiBug

OpenStudy (studygurl14):

8x + 5y = 9 3x + 2y = 4 To solve for a variable, you need to "eliminate" one. You can do that either by multiplying the first equation by -3 and the second by 8, or by multiplyng the first by -2 and the second by 2. (the negatives/positive can be vice versa) I'm not exactly sure what the question is asking for, but I do know how to solve this system of equations

OpenStudy (secret-ninja):

is there anything you can tell me about question B? Do you have any guess what its asking?

OpenStudy (secret-ninja):

8x + 5y = 9 3x + 2y = 4 x=-2 y = 5 Multiplying both equations by 2 16x + 10y = 18 6x +4y =8 x = -2 y = 5

OpenStudy (secret-ninja):

That?

OpenStudy (studygurl14):

Looks correct. Good job. :)